# PQPP: A Joint Benchmark for Text-to-Image Prompt and Query Performance Prediction

### v1

@@ 3.2 Query Performance Assessment
Retrieval models. For text-to-image retrieval, we employ two distinct vision-language models: CLIP [33] and BLIP-2 [24]. For CLIP, we select the base architecture with patches of $32\times 32$ pixels. For BLIP-2, we choose the ViT-Large [16] backbone. Both models are equipped with the necessary components to perform text-to-image retrieval. These models are pre-trained on natural images, which makes them suitable for image retrieval on MS COCO.
@@
Annotation process. For the retrieval setting, we devise a semi-automatic labeling process to generate reference (ground-truth) relevance judgments for the 10,000 image captions chosen as queries. Limited by the annotation budget, we employ an automatic process to restrict the number of retrieved images to 2,000 per query. We exploit the structure of the MS COCO dataset based on (image, caption) pairs to generate preliminary relevance judgments using sentence BERT [42]. More specifically, we compute the cosine similarity in the embedding space of sentence BERT between each query from the established subset and each caption in MS COCO. Based on a preliminary exploratory data analysis, we set the cosine similarity threshold to $0.7$ to determine a comprehensive set of potentially relevant results. The image corresponding to each caption that has the cosine similarity with a query higher than $0.7$ is added to the preliminary set of relevant results for the respective query. For some short and generic text queries, the preliminary set may contain thousands of potentially relevant images. For such queries, we refine the results ranked below 1,000 using a bag-of-words representation. More specifically, a low-rank image is kept only if the bag-of-words representation of the query is included in the bag-of-words representation of the aggregated captions of the respective image. All queries are limited to 2,000 images in the preliminary set. The preliminary steps described above generate a total of 1,245,186 images, which are further subject to rigorous manual review. There are 100 annotators involved in the manual annotation of the potentially relevant images. The annotators are asked to label each image as relevant or irrelevant to the corresponding query. Each image is annotated by two annotators. An image is kept in the relevant set if it is voted as relevant by one annotator. Two of the evaluators, who are also the main authors of this paper, annotated a set of 4,000 queries. These 4,000 queries are used as control queries for the other enrolled annotators. The remaining queries were randomly divided into batches of 50 queries. In each batch, there are 5 control queries, which are used to exclude annotators that provide poor relevance judgments. There are 98 human evaluators who annotated between 1 and 4 batches. We employed the $F_{1}$ measure on relevant images to estimate the quality of the relevance judgments, and set a threshold of $0.4$ to accept annotations. There are 7 annotators who were excluded from the annotation process based on the considered threshold. For the remaining annotators, we obtain a mean $F_{1}$ score of $0.727$. The minimum $F_{1}$ score is $0.447$, which is significantly higher than the $F_{1}$ score of $0.150$ of the random chance baseline. The manual annotation process reduced the total number of relevant images to 431,141. In other words, the annotators removed about a third of the originally retrieved images.
@@
Measuring query performance. To determine the performance level of a query, we employ two alternative measures of retrieval effectiveness, namely the precision for the top $n$ retrieved results (P@$n$) and the reciprocal rank (RR). The precision@$n$ is the ratio between the number of true positive images and $n$. Since P@$10$ is often used in text QPP [48], we adopt the same measure and set $n=10$ for our benchmark. The reciprocal rank of a query is given by the ratio between 1 and the rank of the first relevant result. For example, if the first relevant result is at rank 5, then $RR=1/5=0.2$. We estimate the P@$10$ and RR measures for both CLIP and BLIP-2. Then, the resulting values are averaged across models, thus obtaining an average P@$10$ and an average RR for each query. However, in Appendix A.7, we also report the results per reference system.
@@ 3.3 Evaluation Protocol
We divide the annotated prompts/queries into 6,000 for training, 2,000 for validation, and 2,000 for testing. To evaluate performance predictors, we measure the Pearson and Kendall $\tau$ correlation coefficients between the predicted and the ground-truth performance levels of all test queries, following conventional evaluation procedures in text [48, 51] and image [32] QPP. Furthermore, we test the significance of the results with respect to the random chance baseline using Student’s t-testing [35].
@@ 4.1 Pre-generation/retrieval Predictors
Basic text predictors. Building on linguistic features to predict query difficulty [29], we extract a comprehensive suite of pre-retrieval linguistic indicators: the diversity of concepts (total WordNet synsets per prompt/query), lexical density (number of words per prompt/query), syntactic complexity (average word length), and the frequency of specific grammatical structures (proper nouns, acronyms, numerals, conjunctions and prepositions). We test a wide variety of basic predictors, but we only report the results of the top three predictors, namely the number of synsets, the number of words, and the frequency of conjunctions.
Fine-tuned BERT. We explore the potential of fine-tuning the BERT model [14] as a pre-retrieval performance predictor. We select the base architecture based on cased inputs, since the queries contain named entities. The fine-tuning process involves attaching a custom regression head to the pre-trained BERT backbone, which learns to predict the performance of prompts/queries in image generation and retrieval, respectively. The regression head consists of a dropout layer and two fully connected layers. The dropout rate is set to $0.3$ to prevent overfitting. The first dense layer is based on ReLU activation functions, and it takes the [CLS] token returned by BERT and transforms it into a $512$-dimensional hidden representation. The second layer contains a single neuron (activated by sigmoid) that predicts prompt/query performance. Before training, the ground-truth performance values are normalized to $[0,1]$. We employ grid search on the validation set to establish the optimal hyperparameter configuration. More specifically, we consider learning rates between $10^{-3}$ and $10^{-6}$, and weight decays in the set $\{0,0.1,0.01\}$. All versions are trained for 15 epochs with early stopping, on mini-batches of $256$ samples. We employ AdamW and optimize the mean squared error (MSE) loss. The fine-tuning is independently carried out for the two tasks, generation and retrieval.
@@ 4.2 Post-generation/retrieval Predictors
Fine-tuned CLIP. Our first post-generation/retrieval predictor is based on fine-tuning a CLIP model on (query, image) pairs. We attach a regression head for the image generation scenario, and a binary classification head for the image retrieval scenario. The utilization of CLIP-based embeddings is aimed at leveraging the model’s capacity to jointly represent text and image modalities in a single latent space. The fine-tuned CLIP learns to predict the relevance judgment of each generated or retrieved image for a given prompt/query. The performance of the prompt/query is then predicted by aggregating the predicted relevance judgments. For the generative task, the model uses all four images generated by Stable Diffusion and GLIDE. For the retrieval task, we limit the training data to the first $25$ images returned by each retrieval model. Although $10$ images would be enough for the P@$10$ metric, estimating the RR measure can require more images. A statistical analysis of the training queries indicates that more than 95% of the queries have the first relevant image at a rank higher than $25$, which motivates our choice for limiting the training data to $25$ images per model. The regression/classification head is composed of a two-layer neural network of 512 and 256 neurons, respectively. Both layers are based on ReLU activations. A dropout layer with a drop rate of $0.5$ is added after each dense layer. For the generative task, another layer comprising a single neuron is added to predict prompt performance. The objective of the model is to minimize the MSE loss. For the retrieval task, the last neuron has to determine if an input (query, image) pair is relevant or irrelevant. This is a binary classification task, so the model is trained via binary cross-entropy. We perform a grid search to find the best hyperparameters, considering learning rates between $10^{-3}$ and $10^{-6}$, and weight decays in the set $\{0,0.1,0.01\}$. We employ the AdamW optimizer for $25$ epochs with early stopping, using a batch size of $256$.
[…]
Predictor Generative Task Retrieval Task
Type Predictor Name HBPP P@10 RR
@@
Pre- #synsets $-0.152^{\ddagger}$ $-0.093^{\ddagger}$ $-0.141^{\ddagger}$ $-0.097^{\ddagger}$ $-0.071^{\dagger}$ $-0.037^{\dagger}$
#words $-0.136^{\ddagger}$ $-0.113^{\ddagger}$ $-0.199^{\ddagger}$ $-0.163^{\ddagger}$ $-0.069^{\dagger}$ $-0.052^{\dagger}$
#conjunctions / #words $0.058^{\dagger}$ $0.045^{\dagger}$ $-0.117^{\ddagger}$ $-0.097^{\ddagger}$ $-0.041$ $-0.040$
Fine-tuned BERT $0.568^{\ddagger}$ $0.416^{\ddagger}$ $\textbf{0.507}^{\ddagger}$ $0.318^{\ddagger}$ ${0.262}^{\ddagger}$ ${0.182}^{\ddagger}$
Post- Fine-tuned CLIP $\textbf{0.600}^{\ddagger}$ $\textbf{0.448}^{\ddagger}$ $0.490^{\ddagger}$ $\textbf{0.330}^{\ddagger}$ ${0.197}^{\ddagger}$ ${0.129}^{\ddagger}$
Correlation CNN $0.471^{\ddagger}$ $0.347^{\ddagger}$ $0.327^{\ddagger}$ $0.196^{\ddagger}$ $\textbf{0.301}^{\ddagger}$ $\textbf{0.212}^{\ddagger}$
Correlation-based CNN. Following the work of Sun et al. [41], we employ a convolutional neural network (CNN) trained on the set of generated or retrieved images, respectively. The CNN model takes a correlation matrix between all image pairs as input. For the generative task, the size of the input correlation matrix is $4\times 4$. For the retrieval task, we apply the same limit to the number of retrieved images per query as for the fine-tuned CLIP predictor. Hence, the size of the correlation matrix for one retrieval model is $25\times 25$. We concatenate the correlation matrices for CLIP and BLIP-2 models in the channel dimension, which results in a tensor of $25\times 25\times 2$ components, that is given as input to the CNN. The correlation of an (image, image) pair is given by the cosine similarity between the (pre-trained) CLIP embeddings of the respective images. The CNN architecture is composed of four convolutional-pooling blocks, followed by two linear layers. This is a custom architecture that comprises $3\times 3$ convolutional filters applied at a stride of $1$, using a padding of $1$. The number of filters in each of the four convolutional layers is 64, 128, 256 and 512, respectively. A max-pooling is applied after each convolutional layer. The pooling operation uses $2\times 2$ filters applied at a stride of $2$. The first fully connected layer comprises 1024 units. Each hidden neuron is followed by a ReLU activation. The final layer comprises a single neuron that is trained in a regression setting via the MSE loss. The hyperparameter tuning is identical to the one employed for the fine-tuned CLIP model. The correlation-based CNN is trained for $25$ epochs using AdamW with early stopping, on mini-batches of $256$ samples.
5 Experiments and Results
Figure 3: t-SNE visualization of the test prompts embedded in the latent space of the fine-tuned BERT on the generation task. The ground-truth HBPP performance is encoded via a color map from green to red. The visualization confirms that the fine-tuned BERT predictor learns a meaningful representation of the prompts. Best viewed in color.
In Table 2, we report the results of the established predictors for both generative and retrieval tasks, where the reference or ground truth is the average of the two reference systems. We provide the separate results for each reference system in Appendix A.7. For the generative task, the highest correlation between the predicted and the actual performance corresponds to the fine-tuned CLIP, with a Pearson correlation coefficient of $0.6$. This predictor shows a lower correlation with P@$10$ in the retrieval task, where the Pearson score is $0.49$. In general, predictors exhibit higher correlations with HBPP than with P@$10$. The best predictor among the pre-generation/retrieval ones is the fine-tuned BERT, for both tasks and all evaluation measures. Its correlation measures clearly indicate a positive correlation, e.g. the Pearson correlation with P@$10$ on the retrieval task is as high as $0.507$. The two post-generation/retrieval predictors exhibit good performance levels. The fine-tuned CLIP provides better results for the HBPP and P@$10$ measures, while the correlation CNN is more suitable for the RR measure. Overall, an interesting finding of our study is that the pre-retrieval fine-tuned BERT is a worthy competitor for the better post-retrieval predictor (fine-tuned CLIP). In Table 2, we only report the predictors that obtain a correlation higher than $0.1$, in at least one evaluation scenario. The results of the full set of predictors are presented in Appendix A.7. The very limited effectiveness of the excluded pre-generation/retrieval predictors can be attributed to their behavior of capturing only a limited portion of the relevant information (see also Appendix A.6).
In Figure 3, we present a t-SNE visualization of the test prompts embedded in the latent space of the BERT predictor fine-tuned on the generation task. We observe that the learned embedding correlates well with the ground-truth HBPP values, explaining the high accuracy of the fine-tuned BERT predictors on the generation task. An analogous visualization for the retrieval task is shown in Figure 6 from Appendix A.8.
Additional results. In Appendix A.6, we discuss empirical evidence regarding the overlap between generation and retrieval performance. In Appendix A.7, we present the complete set of quantitative results. In Appendix A.8, we analyze the generation/retrieval results from a qualitative perspective.
6 Conclusion and Future Work
Benchmark collections lay the grounds for the development of effective approaches to solve new challenges. In this paper, we have developed the first manually-labeled benchmark for prompt performance prediction in the context of prompt-to-image generation. Our benchmark is also applicable for query-to-image retrieval, enabling the direct comparison of the performance prediction task in generation vs. retrieval scenarios. Our dataset and code are made publicly available to the research community.
One direction for future work is to develop a model that combines different pre-retrieval predictors. Combining predictors in a supervised manner has shown its effectiveness in text [13] and image [32] QPP. Another important direction is to organize a shared task associated with the proposed PQPP benchmark, so that the IR community can further explore the novel task of prompt performance prediction in image generation.
Appendix A Appendix
A.1 Limitations
In the retrieval setting, the initial ground-truth image bank was generated using a method that combines Sentence-BERT and the bag-of-words model. We acknowledge that some images may contain content that is not fully or accurately captured by their paired captions, which could introduce occasional false negatives or false positives in the ground-truth collection. This heuristic was chosen to minimize the number of false negatives, while the false positives were mostly curated by the enrolled annotators.
Another limitation of our study is that we consider prompts that correspond to naturally occurring images. However, generative models are often used to generate fictive images. Since retrieval models can only retrieve images from a database of real images, queries that correspond to fictive images can be easily labeled as difficult. We therefore choose captions from MS COCO to serve as prompts/queries for our benchmark, creating an evaluation setting that is representative for both generation and retrieval.
A.2 Potential Negative Societal Impact
The development and deployment of text-to-image generation and retrieval systems come with several societal implications that warrant careful consideration. Here, we outline key areas of concern. Our enhanced dataset is built upon a preexisting large dataset, which may inherit and perpetuate biases present in the original data. Additionally, user annotations might have been influenced by their own cultural backgrounds, potentially introducing subjective biases into the final decisions. The pre-trained generative models employed in our study could also exhibit inherent biases, affecting the generated outputs. These combined factors could lead to unfair representations or reinforce existing stereotypes. We acknowledge the necessity of ongoing efforts to identify, measure, and mitigate these biases to ensure the fairness and inclusiveness of our models. In our work, we address these concerns by promoting transparency in our methodology and unifying multiple user annotations to mitigate possible individual biases.
A.3 Ethical Considerations on Data Annotation
Data annotation by students is a common practice in our three universities (University of Bucharest, Aix-Marseille Université, Université Toulouse Jean-Jaurès) and we followed the standard protocols to get approvals from the corresponding ethics committees. The enrolled students were compensated with bonus points. We would like to emphasize that the students understood that the annotation task is optional, and they could also get the extra bonus points by performing alternative tasks (which did not involve data annotation). Moreover, all students were given the opportunity to obtain a full grade without the optional annotation task. Hence, there was no obligation for any of the students to perform the annotations. The students were also able to opt out, at any time during the annotation, without any penalties.
The enrolled students come from three universities from two different countries (France and Romania). They represent different cultural, gender and age groups, reflecting the varied distribution of students in each of the three universities.
A.4 Related Work on QPP in Text Retrieval
Pre-retrieval features are based on information available prior to the execution of the query. Some are independent to the document collection, such as query length, part-of-speech features (including the number of words of some grammatical categories), query ambiguity [12], and query complexity [29]. Other pre-retrieval features depend on the document collection statistics, such as the inverse document frequency [40], the query scope [22] (which measures the coverage of a query within the context of a document collection and estimates the proportion of documents that are relevant to the query), and the SQC [51] (which is a similarity score between the query and the collection). Pre-retrieval predictors have the huge advantage of being determined before running the search, but they have been found to be less effective than post-retrieval ones on textual ad hoc retrieval [20, 29, 34, 38]. Unlike pre-retrieval features, post-retrieval features require conducting document retrieval with the query. Most of these features are calculated based on the scores of the retrieved documents, quantifying the robustness of the document list, or considering the distribution of the document scores [6, 8, 9, 34, 50]. On textual IR, the Clarity Score estimates the specificity of a query considering the language distribution of the document collection and that of the top-retrieved documents [8]. The Normalized Query Commitment (NQC), also known as the query drift [38], measures how much the retrieved documents deviate from the central topic of the query. The Weighted Information Gain (WIG) calculates the difference in information content between the documents retrieved for a specific query and a baseline distribution of information in the collection or corpus, based on the scores of the top-retrieved documents [52].
More recent work investigated QPP on neural IR (NIR) systems [2, 10, 18, 27, 39, 49]. Datta et al. [10] employed convolutional neural layers for their Deep-QPP predictor. This architecture has further been combined with LETOR post-retrieval predictors with some success [11]. According to Faggioli et al. [18], QPP models which have been developed for sparse IR methods perform worse when applied to NIR systems. However, the authors did not consider linguistic-based predictors in their work. On the other hand, supervised BERT-based QPP models seem to work better. Arabzadeh et al. [3] used BERT to predict the performance of search queries in terms of their ability to retrieve relevant documents from a corpus. Such predictors may better capture the semantic aspects of the query-document matching.
A.5 Distribution of Human Relevance Judgments for Prompt Performance Prediction
A.6 Generation vs. Retrieval Performance
[…]
Metric Pearson Kendall
HBPP vs. P@10 0.118 0.082
HBPP vs. RR 0.070 0.046
P@10 vs. RR 0.562 0.504
A.7 Detailed Quantitative Results
[…]
Predictor Generative Task
Type Predictor Name Stable Diffusion GLIDE
Pearson
Kendall
Pearson
Kendall
Pre- #synsets $-0.115^{\ddagger}$ $-0.085^{\ddagger}$ $-0.123^{\ddagger}$ $-0.082^{\ddagger}$
#words $-0.118^{\ddagger}$ $-0.115^{\ddagger}$ $-0.101^{\ddagger}$ $-0.092^{\ddagger}$
#conjunctions / #words $0.037$ $-0.011$ $0.051$ $0.041$
Fine-tuned BERT $\textbf{0.252}^{\ddagger}$ $\textbf{0.149}^{\ddagger}$ $0.558^{\ddagger}$ $0.405^{\ddagger}$
Post- Fine-tuned CLIP $0.247^{\ddagger}$ $0.141^{\ddagger}$ $\textbf{0.615}^{\ddagger}$ $\textbf{0.461}^{\ddagger}$
Correlation CNN $0.158^{\ddagger}$ $0.061^{\ddagger}$ $0.505^{\ddagger}$ $0.368^{\ddagger}$
[…]
Predictor Retrieval Task (CLIP)
Type Predictor Name P@10 RR
Pearson
Kendall
Pearson
Kendall
Pre- #synsets $-0.129^{\ddagger}$ $-0.080^{\ddagger}$ $-0.052$ $-0.035$
#words $-0.161^{\ddagger}$ $-0.132^{\ddagger}$ $-0.054$ $-0.053$
#conjunctions / #words $-0.086^{\ddagger}$ $-0.073^{\ddagger}$ $-0.029$ $-0.029$
Fine-tuned BERT ${0.441}^{\ddagger}$ $0.272^{\ddagger}$ $0.218^{\ddagger}$ $0.157^{\ddagger}$
Post- Fine-tuned CLIP $\textbf{0.458}^{\ddagger}$ $\textbf{0.299}^{\ddagger}$ ${0.173}^{\ddagger}$ ${0.120}^{\ddagger}$
Correlation CNN $0.314^{\ddagger}$ $0.180^{\ddagger}$ $\textbf{0.279}^{\ddagger}$ $\textbf{0.210}^{\ddagger}$
[…]
Predictor Retrieval Task (BLIP-2)
Type Predictor Name P@10 RR
Pearson
Kendall
Pearson
Kendall
[…]
#words $-0.206^{\ddagger}$ $-0.167^{\ddagger}$ $0.058^{\dagger}$ $0.042^{\ddagger}$
#conjunctions / #words $-0.128^{\ddagger}$ $-0.108^{\ddagger}$ $-0.038$ $-0.039$
Fine-tuned BERT $\textbf{0.499}^{\ddagger}$ $0.317^{\ddagger}$ $0.210^{\ddagger}$ $0.157^{\ddagger}$
Post- Fine-tuned CLIP $0.457^{\ddagger}$ $\textbf{0.339}^{\ddagger}$ ${0.149}^{\ddagger}$ ${0.115}^{\ddagger}$
Correlation CNN $0.298^{\ddagger}$ $0.190^{\ddagger}$ $\textbf{0.213}^{\ddagger}$ $\textbf{0.160}^{\ddagger}$
[…]
Predictor Generative Task Retrieval Task
Type Predictor Name HBPP P@10 RR
@@ A.7 Detailed Quantitative Results
Pre- #synsets $-0.152^{\ddagger}$ $-0.093^{\ddagger}$ $-0.141^{\ddagger}$ $-0.097^{\ddagger}$ $-0.071^{\dagger}$ $-0.037^{\dagger}$
#words $-0.136^{\ddagger}$ $-0.113^{\ddagger}$ $-0.199^{\ddagger}$ $-0.163^{\ddagger}$ $-0.069^{\dagger}$ $-0.052^{\dagger}$
avg word length $0.049$ $0.039^{\dagger}$ $-0.116^{\ddagger}$ $-0.084^{\ddagger}$ $-0.067^{\dagger}$ $-0.046^{\dagger}$
avg #acronyms $0.010$ $0.004$ $-0.001$ $-0.021$ $0.015$ $0.012$
avg #numerals $-0.045$ $-0.036$ $-0.065^{\dagger}$ $-0.066^{\ddagger}$ $-0.018$ $-0.001$
#conjunctions / #words $0.058^{\dagger}$ $0.045^{\dagger}$ $-0.117^{\ddagger}$ $-0.097^{\ddagger}$ $-0.041$ $-0.040$
#prepositions / #words $0.032$ $0.013$ $0.060^{\dagger}$ $0.034$ $0.035$ $0.024$
edge count $0.037$ $0.062^{\ddagger}$ $0.027$ $0.047$ $0.013$ $0.010$
edge weight sum $0.030$ $0.057^{\ddagger}$ $0.025$ $0.046^{\dagger}$ $0.018$ $0.004$
inverse weight freq $0.090^{\ddagger}$ $0.071^{\ddagger}$ $0.045$ $0.039$ $0.010$ $0.011$
degree centrality $0.027$ $0.005$ $0.008$ $0.030$ $0.028$ $0.001$
cloneness centrality $0.085^{\ddagger}$ $0.065^{\ddagger}$ $0.048$ $0.038$ $0.016$ $0.011$
between centrality $0.020$ $0.004$ $0.007$ $0.016$ $0.020$ $0.004$
Page Rank $0.086$ $0.064$ $0.049$ $0.040^{\dagger}$ $0.014$ $0.011$
Fine-tuned BERT $0.568^{\ddagger}$ $0.416^{\ddagger}$ $\textbf{0.507}^{\ddagger}$ $0.318^{\ddagger}$ $0.262^{\ddagger}$ $0.182^{\ddagger}$
Post- Fine-tuned CLIP $\textbf{0.600}^{\ddagger}$ $\textbf{0.448}^{\ddagger}$ $0.490^{\ddagger}$ $\textbf{0.330}^{\ddagger}$ ${0.197}^{\ddagger}$ ${0.129}^{\ddagger}$
Correlation CNN $0.471^{\ddagger}$ $0.347^{\ddagger}$ $0.327^{\ddagger}$ $0.196^{\ddagger}$ $\textbf{0.301}^{\ddagger}$ $\textbf{0.212}^{\ddagger}$
Score Variance @100 $0.003$ $0.005$ $0.016$ $0.003$ $0.016$ $0.004$
In Table 4, we present the results of prompt performance predictors for each generative model (Stable Diffusion and GLIDE). We observe that predicting prompt performance for Stable Diffusion is generally more difficult than for GLIDE. This observation can be attributed to the higher variety (in terms of relevance to the prompt) of the images generated by GLIDE.
In Table 5 and Table 6, we report the results of query performance predictors for the CLIP and BLIP-2 retrieval systems, respectively. The results are consistent with the ones presented in Table 2, which are obtained by averaging the performance of the two models as a ground truth. Indeed, the pre-retrieval predictors based on the number of synsets, the number of words, and the ratio between conjunctions and words are weakly correlated with the individual system performance. For the RR performance measure, both Pearson and Kendall correlations indicate that the best predictor for both retrieval models is the correlation CNN. For P@10, the correlation for the pre-retrieval fine-tuned BERT and the post-retrieval fine-tuned CLIP are very close to each other, for both retrieval models and both correlation measures.
In Table 7, we present the results for all the considered predictors, although as mentioned in Section 5, Table 2 shows the most interesting predictors. It is nevertheless important to also report failed attempts on specific predictors. In general, we find that neural predictors clearly outperform predictors that are based on simple heuristics.
A.8 Qualitative Results

### v2

@@ 3.2 Query Performance Assessment
Retrieval models. For text-to-image retrieval, we employ two distinct vision-language models: CLIP [37] and BLIP-2 [27]. For CLIP, we select the ViT-Base [19] architecture with patches of $32\times 32$ pixels. For BLIP-2, we choose the ViT-Large backbone. These models are pre-trained on natural images, which makes them suitable for image retrieval on MS COCO.
@@
Annotation process. For the retrieval setting, we devise a semi-automatic labeling process to generate reference (ground-truth) relevance judgments for the 10,200 queries. We first employ an automatic process to restrict the number of retrieved images to 2,000 per query. We exploit the structure of the MS COCO dataset based on (image, caption) pairs to generate preliminary relevance judgments using sentence BERT [47]. More specifically, we compute the cosine similarity in the embedding space of sentence BERT between each query in PQPP and each caption in MS COCO. Based on a preliminary exploratory data analysis, we set the cosine similarity threshold to $0.7$ to determine a comprehensive set of potentially relevant results. The image corresponding to each caption that has the cosine similarity with a query higher than $0.7$ is added to the preliminary set of relevant results for the respective query. For some short and generic text queries, the preliminary set may contain thousands of potentially relevant images. For such queries, we refine the results ranked below 1,000 using a bag-of-words representation. More specifically, a low-rank image is kept only if the bag-of-words representation of the query is included in the bag-of-words representation of the aggregated captions of the respective image. All queries are limited to 2,000 images in the preliminary set that undergoes manual labeling. The preliminary steps described above generate a total of 1,393,363 images, which are further subject to rigorous manual review. There are 100 annotators involved in the manual annotation of the potentially relevant images. The annotators are asked to label each image as relevant or irrelevant to the corresponding query. Each image is annotated by two annotators. An image is kept in the relevant set if it is voted as relevant by one annotator. Two of the evaluators, who are also the main authors of this paper, annotated a set of 4,200 queries. These 4,200 queries are used as control queries for the other enrolled annotators. The remaining queries were randomly divided into batches of 50 queries. In each batch, there are 5 control queries, which are used to exclude annotators that provide poor relevance judgments. There are 98 human evaluators who annotated between 1 and 4 batches. We employed the $F_{1}$ measure on relevant images for control queries to estimate the quality of the relevance judgments, and set a threshold of $0.4$ to accept annotations. There are 7 annotators who were excluded from the annotation process based on the considered threshold. For the remaining annotators, we obtain a mean $F_{1}$ score of $0.727$. The minimum $F_{1}$ score is $0.447$, which is significantly higher than the $F_{1}$ score of $0.150$ of the random chance baseline. The manual annotation process reduced the total number of relevant images to 530,360. In other words, the annotators removed almost two thirds of the originally retrieved images.
@@
Measuring query performance. To determine the performance level of system for a given query, we employ two alternative measures of retrieval effectiveness, namely the precision for the top $n$ retrieved results (P@$n$) and the reciprocal rank (RR). The precision@$n$ is the ratio between the number of true positive images and $n$. Since P@$10$ is often used in text QPP [56], we adopt the same measure and set $n=10$ for our benchmark. The reciprocal rank of a query is given by the ratio between 1 and the rank of the first relevant result. We estimate the P@$10$ and RR measures for both CLIP and BLIP-2.
@@ 3.3 Evaluation Protocol
We divide the annotated prompts/queries into 6,080 for training, 2,040 for validation, and 2,080 for testing. To evaluate performance predictors, we measure the Pearson and Kendall $\tau$ correlation coefficients between the predicted and the ground-truth performance levels of all test queries, following conventional evaluation procedures in text [56, 60] and image [36] QPP. Furthermore, we test the significance of the results with respect to the random chance baseline using Student’s t-testing [39].
4 Generation vs. Retrieval Performance
Metric Pearson Kendall
HBPP vs. P@10 0.135 0.093
HBPP vs. RR 0.072 0.048
P@10 vs. RR 0.560 0.512
[…]
@@ 5 Predictors
We briefly present the chosen performance predictors below. We provide details about implementation choices and hyperparameter tuning in Appendix 11.
@@ 5.1 Pre-generation/retrieval Predictors
Basic text predictors. Building on linguistic features to predict query difficulty [32], we extract a comprehensive suite of pre-retrieval linguistic indicators: the diversity of concepts (total WordNet synsets per prompt/query), lexical density (number of words per prompt/query), morphological complexity (average word length), and the frequency of specific grammatical structures (proper nouns, acronyms, numerals, conjunctions and prepositions). We test a wide variety of basic predictors, but we only report the results of the top two predictors, namely the number of synsets (#synsets) and the number of words (#words). Results with the other basic text predictors are reported in Appendix 12.
Fine-tuned BERT. We explore the potential of fine-tuning the BERT model [17] as a pre-retrieval performance predictor. We select the base architecture based on cased inputs, since the queries contain named entities. The fine-tuning process involves attaching a custom regression head to the pre-trained BERT backbone, which learns to predict the performance of prompts/queries in image generation and retrieval, respectively.
@@ 5.2 Post-generation/retrieval Predictors
Fine-tuned CLIP. Our first post-generation/retrieval predictor is based on fine-tuning a CLIP model on (query, image) pairs. We use Long-CLIP [58] with a ViT-B/32 backbone. We attach a regression head for the image generation scenario, and a binary classification head for the image retrieval scenario. The utilization of CLIP-based embeddings is aimed at leveraging the model’s capacity to jointly represent text and image modalities in a single latent space. The fine-tuned CLIP learns to predict the relevance judgment of each generated or retrieved image for a given prompt/query. The performance of the prompt/query is then predicted by aggregating the predicted relevance judgments.
Predictor Type Predictor Name Generative Task Retrieval Task
GLIDE SDXL CLIP BLIP-2
HBPP HBPP P@10 RR P@10 RR
@@
Pre- #synsets $-0.112^{\ddagger}$ $-0.076^{\ddagger}$ $-0.087^{\ddagger}$ $-0.080^{\ddagger}$ $-0.110^{\dagger}$ $-0.058^{\ddagger}$ $-0.034$ $-0.012$ $-0.115^{\ddagger}$ $-0.070^{\ddagger}$ $-0.038$ $-0.010$
#words $-0.090^{\dagger}$ $-0.084^{\ddagger}$ $-0.105^{\ddagger}$ $-0.109^{\ddagger}$ $-0.133^{\ddagger}$ $-0.104^{\ddagger}$ $-0.035$ $-0.026$ $-0.175^{\ddagger}$ $-0.136^{\ddagger}$ $-0.038$ $-0.015$
Fine-tuned BERT $0.566^{\ddagger}$ $0.406^{\ddagger}$ $0.281^{\ddagger}$ $0.232^{\ddagger}$ $0.451^{\ddagger}$ $0.277^{\ddagger}$ $\mathbf{0.221}^{\ddagger}$ $\mathbf{0.176}^{\ddagger}$ $\mathbf{0.511}^{\ddagger}$ $0.328^{\ddagger}$ $0.168^{\ddagger}$ $0.139^{\ddagger}$
Post- Fine-tuned CLIP $\mathbf{0.649}^{\ddagger}$ $\mathbf{0.474}^{\ddagger}$ $\mathbf{0.380}^{\ddagger}$ $\mathbf{0.246}^{\ddagger}$ $\mathbf{0.473}^{\ddagger}$ $\mathbf{0.299}^{\ddagger}$ $0.200^{\ddagger}$ $0.149^{\ddagger}$ $0.498^{\ddagger}$ $\mathbf{0.358}^{\ddagger}$ $0.166^{\ddagger}$ $0.150^{\ddagger}$
Correlation CNN $0.548^{\ddagger}$ $0.393^{\ddagger}$ $0.159^{\ddagger}$ $0.107^{\ddagger}$ $0.270^{\ddagger}$ $0.186^{\ddagger}$ $0.189^{\ddagger}$ $0.162^{\ddagger}$ $0.159^{\ddagger}$ $0.133^{\ddagger}$ $\mathbf{0.206}^{\ddagger}$ $\mathbf{0.158}^{\ddagger}$
[…]
Predictor Type Predictor Name Generative Task Retrieval Task
[…]
HBPP HBPP P@10 RR P@10 RR
Pearson
Kendall
Pearson
Kendall
Pearson
Kendall
Pearson
Kendall
Pearson
Kendall
Pearson
Kendall
Pre- Fine-tuned BERT $0.165^{\ddagger}$ $0.128^{\ddagger}$ $0.150^{\ddagger}$ $0.087^{\dagger}$ $0.420^{\ddagger}$ $0.264^{\ddagger}$ $0.134^{\dagger}$ $0.112^{\ddagger}$ $0.431^{\ddagger}$ $0.270^{\ddagger}$ $\mathbf{0.165}^{\ddagger}$ $0.133^{\ddagger}$
Post- Fine-tuned CLIP $\mathbf{0.256}^{\ddagger}$ $\mathbf{0.186}^{\ddagger}$ $\mathbf{0.179}^{\ddagger}$ $\mathbf{0.089}^{\dagger}$ $\mathbf{0.449}^{\ddagger}$ $\mathbf{0.291}^{\ddagger}$ $0.121^{\dagger}$ $0.122^{\ddagger}$ $\mathbf{0.453}^{\ddagger}$ $\mathbf{0.337}^{\ddagger}$ $0.151^{\ddagger}$ $\mathbf{0.154}^{\ddagger}$
Correlation CNN $0.131^{\ddagger}$ $0.078^{\ddagger}$ $0.096^{\dagger}$ $0.020$ $0.155^{\ddagger}$ $0.130^{\ddagger}$ $\mathbf{0.206}^{\ddagger}$ $\mathbf{0.162}^{\ddagger}$ $0.228^{\ddagger}$ $0.167^{\ddagger}$ $0.155^{\ddagger}$ $0.130^{\ddagger}$
[…]
Correlation-based CNN. Following the work of Sun et al. [46], we employ a convolutional neural network (CNN) trained on the set of generated or retrieved images, respectively. The CNN model takes a correlation matrix between all image pairs as input. The correlation of an (image, image) pair is given by the cosine similarity between the (pre-trained) CLIP embeddings of the respective images.
6 Experiments and Results